A diver comes off a board with arms straight up and legs straight down, giving her a moment of inertia about her rotation axis of . She then tucks into a small ball, decreasing this moment of inertia to . While tucked, she makes two complete revolutions in 1.3 .


If she hadn't tucked at all, how many revolutions would she have made in the 1.7 from board to water?
Express your answer using two significant figures.

=_____________rev
Answer #1
Concepts and reason

The concept required to solve the given problem is law of conservation of angular momentum.

Initially, find the angular speed while she tucked i.e. ω2{\omega _2} . Use the law of conservation of angular momentum, re-arrange that for ω1{\omega _1} , substitute the values in the expression and find ω1{\omega _1} . Find the number of revolutions made by her in 1.7 sec.

Fundamentals

The law of conservation of angular momentum states that the angular momentum of a body is conserved unless an external torque acts on the system.

Li=Lf{L_i} = {L_f}

Here, i and f represent initial and final states of a body.

The moment of inertial value in the straight position is,

I1=18kg/m2{I_1} = 18{\rm{ kg/}}{{\rm{m}}^2}

The moment of inertial value when she tucks is,

I2=3.6kg/m2{I_2} = 3.6{\rm{ kg/}}{{\rm{m}}^2}

Angular speed while she tucked is calculated as:

ω2=Nt{\omega _2} = \frac{N}{t}

Here, N is the number of revolutions and t is the time taken by her to complete that revolutions.

Substitute 2 rev for N and 1.3 sec for t in equation ω2=Nt{\omega _2} = \frac{N}{t} .

ω2=2rev1.3sec=1.54rev/sec\begin{array}{c}\\{\omega _2} = \frac{{2{\rm{ rev}}}}{{1.3{\rm{ sec}}}}\\\\ = 1.54{\rm{ rev/sec}}\\\end{array}

The law of conservation of angular momentum states that the angular momentum of a body is conserved unless an external torque acts on the system.

Li=Lf{L_i} = {L_f}

The above equation can be reduced as:

I1ω1=I2ω2{I_1}{\omega _1} = {I_2}{\omega _2}

Here, I1{I_1} and I2{I_2} are the initial and final moment of inertia and ω1{\omega _1} and ω2{\omega _2} are the initial and final angular speed of the system.

Re-arrange equation I1ω1=I2ω2{I_1}{\omega _1} = {I_2}{\omega _2} for ω1{\omega _1} .

ω1=(I2I1)ω2{\omega _1} = \left( {\frac{{{I_2}}}{{{I_1}}}} \right){\omega _2}

Substitute 1.54 rev/sec for ω2{\omega _2} , 18kg/m218{\rm{ kg/}}{{\rm{m}}^2} for I1{I_1} and 3.6kg/m23.6{\rm{ kg/}}{{\rm{m}}^2} for I2{I_2} in equation ω1=(I2I1)ω2{\omega _1} = \left( {\frac{{{I_2}}}{{{I_1}}}} \right){\omega _2} .

ω1=(3.6kg/m218kg/m2)(1.54rev/sec)=0.308rev/sec\begin{array}{c}\\{\omega _1} = \left( {\frac{{3.6{\rm{ kg/}}{{\rm{m}}^2}}}{{18{\rm{ kg/}}{{\rm{m}}^2}}}} \right)\left( {1.54{\rm{ rev/sec}}} \right)\\\\ = 0.308{\rm{ rev/sec}}\\\end{array}

Number of revolutions made in 1.7 sec are,

N=ω1(1.7sec)N = {\omega _1}\left( {1.7{\rm{ sec}}} \right)

Substitute 0.308 rev/sec for ω1{\omega _1} in N=ω1(1.7sec)N = {\omega _1}\left( {1.7{\rm{ sec}}} \right) .

N=(0.308rev/sec)(1.7sec)=0.52rev\begin{array}{c}\\N = \left( {0.308{\rm{ rev/sec}}} \right)\left( {1.7{\rm{ sec}}} \right)\\\\ = 0.52{\rm{ rev}}\\\end{array}

Ans:

The number of revolutions is 0.52.